# Taming Diffusion Transformer for Real-Time Mobile Video Generation

### v1

@@ 6 Conclusion, Limitations and Broader Impact
Despite these advances, our method has several limitations. First, the highly compressed latent space and DiT pruning lead to occasional degradations in fine-grained details, particularly in fast motion or complex texture scenes. Second, due to various practical constraints, most state-of-the-art video diffusion models (VDMs) used for comparison in this work, including our own, are trained on internally collected video datasets that cannot be fully disclosed or released. As a result, direct comparisons may not be entirely fair and reproducible. To mitigate this limitation, we include a reproduction of the LTX model trained on our dataset and report the results in the supplementary material. This work enables efficient video generation on mobile devices, but also carries the potential risk of misuse for generating fake or inappropriate content.

### v2

@@ Appendix I Impact of Inference Steps
Despite these advances, our method has several limitations. First, the highly compressed latent space and DiT pruning lead to occasional degradations in fine-grained details, particularly in fast motion or complex texture scenes. Second, due to various practical constraints, most state-of-the-art video diffusion models (VDMs) used for comparison in this work, including our own, are trained on internally collected video datasets that cannot be fully disclosed or released. As a result, direct comparisons may not be entirely fair and reproducible. To mitigate this limitation, we include a reproduction of the LTX model trained on our dataset and report the results in the Appendix˜J. This work enables efficient video generation on mobile devices, but also carries the potential risk of misuse for generating fake or inappropriate content.
Appendix O More Qualitative Results